OpenStudy (anonymous):

When a mole of lithium reacts chemically with hydrogen to form lithium hydride, the energy released is found to be roughly 90 joules. How much mass is lost? [(1 J = 1 kg times m squared over s squared ) and the speed of light = 3.00 × 108 m over s] Answer 1.0 × 10-15 kg 2.7 × 1010 kg 1.0 × 1015 kg 3.0 × 10-7 kg

OpenStudy (abhisar):

Let the mass lost be m, then using \(\sf E=mc^2\) \(\sf 90=m\times (3\times 10^8)^2\) Now this equation for m

OpenStudy (abhisar):

*Now solve this equation for m

OpenStudy (anonymous):

how do i work that problem?

OpenStudy (abhisar):

??? Do u want me to tell you how did i set up that equation or u need help in solving that equation ?

OpenStudy (anonymous):

sorry i really am sleep deprived, but i need help solving :/

OpenStudy (abhisar):

Then u should get some sleep :p

OpenStudy (anonymous):

to much to do and to much going on with legal work to sleep. probably couldnt if i tried anyway

OpenStudy (abhisar):

\(\sf 90=m\times (3\times 10^8)^2\\ \Rightarrow 90=m\times 9\times 10^{16}\\ \Rightarrow m=\LARGE \frac{90}{9\times 10^{16}}\\ m=?\)

OpenStudy (anonymous):

1 x 10 to the 15?

OpenStudy (abhisar):

\(\sf 1\times 10^{-15}\)

OpenStudy (anonymous):

always miss the negative

OpenStudy (abhisar):

is it clear ?

OpenStudy (abhisar):

@gstprc

OpenStudy (anonymous):

yes thanks again. online classes are harder to learn by especially when its a course like chemistry

OpenStudy (abhisar):

Did you know you can give medals to people who help you with questions? Just click the \(\color{blue}{\text{blue "best response" button}}\) to show appreciation for the people who help you.

OpenStudy (abhisar):

\(\rlap{\color{blue}{\huge\bigstar}\huge \color{blue}{ \text{You're Most Welcome! }}\color{blue}\bigstar}{\; \color{aqua}{\huge\bigstar}\huge \color{aqua}{\text{You're Most Welcome! }}\color{aqua}\bigstar}\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)

OpenStudy (anonymous):

you realize i have been giving you the medals on all these questions lol